OpenStudy (anonymous):

Given the exponential function f(x) = 5(2)^x, evaluate f(-2)

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

sub the x value of -2 into the equation gives 5/4 ok?

OpenStudy (anonymous):

Look, 2 ^ -2 = 1/(2)^2 = 1/4 Multiplying by 5, Answer is 5/4. Understood?

OpenStudy (anonymous):

Thanks!!!I understand it now :)